A wildlife biologist caught, banded, and released 400 quail at a local quail ranch. One month later, he captured 100 quail in the same reserve and discovered that 16 of these were banded. Estimate the number of quail on the quail ranch and match your result to the correct answer below.
Answer by checkley77(9273)

16/100=400/x
16x=100*400
16x=40,000
x=40,000/16
x=2,500 answer for the projected number of quail.